OpenStudy (iloveleatiforever):

Solve for d. 3d + 1 = 10

OpenStudy (andrejrusa12):

D=3 3*3=9+1=10

OpenStudy (mathmale):

You need to isolate "d" on the left side of your equation. To accomplish this, please subtract 1 from both sides of the equation. What do you get as a result?

OpenStudy (mathstudent55):

You want d alone. d is being multiplied by 3, and then 1 is being added to the product. You need to undo both the addition and the multiplication. If you knew what d is, you'd multiply d by 3, then you'd add 1. To undo the operations, you do the opposite operations in the reverse order. Start by subtracting 1 from both sides. Then divide both sides by 3.